OpenStudy (anonymous):

Using complete sentences, explain how to graph 2x + y = 14 using the x- and y-intercepts method.

OpenStudy (anonymous):

Are you familiar with the equation y=mx+b ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Well my first hint would be to describe how you would make that equation look like y=mx+b So solve for y and then post back your answer and i'll let you know if you're correct. we can go from there.

OpenStudy (anonymous):

i did it

OpenStudy (anonymous):

but i get 22/98

OpenStudy (anonymous):

Can you what I am saying @Diana97

OpenStudy (anonymous):

*do missing..

OpenStudy (anonymous):

Oh no no. Solve the equation so you get y= something. I'll give you another hint. subtract 2x over.

OpenStudy (anonymous):

@waterineyes wat

OpenStudy (anonymous):

@agentc0re can you give me 5 minutes so that I can explain??

OpenStudy (anonymous):

See you have to find using x and y intercept method: For finding x intercept Put y = 0 in the given equation can you do that?? And tell me what you got for x...?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

you have 2x + y = 14 Here put y = 0 and find x..

OpenStudy (anonymous):

is x

OpenStudy (anonymous):

No do it again and be careful..

OpenStudy (anonymous):

Simply put y = 0..

OpenStudy (anonymous):

Are you getting ?? If not getting then tell me fastly don't look here and there..

OpenStudy (anonymous):

@Diana97 Are you listening to me or not???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

im doin it

OpenStudy (anonymous):

Then find what I have told you.. Take your time and tell me the value of x after taking that much of your time..

OpenStudy (anonymous):

x represent 2'

OpenStudy (anonymous):

How can you show me what you did...??

OpenStudy (anonymous):

this wat i did First set x equal to zero so y equals to 14. Then set y equal to zero so 2x equals 14. Divide the 2 by 14 and 2x so you get x by itself. You are left with x equals to 7. Your coordinates are then (0,14) and (7,0).

OpenStudy (anonymous):

Then what are you doing above???

OpenStudy (anonymous):

You have found that absolutely correct..

OpenStudy (anonymous):

Now can you graph them or not???

OpenStudy (anonymous):

no i can't

OpenStudy (anonymous):

 Getting how to graph??